20. If 7 times the 7th term of an A.P. is equal to 11 times its 11th term, then 18th term is
(a) 18
(b) 9
(c) 77
(d) 0

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: d Explaination:Reason: We have 7a7 = 11a11 ⇒ 7[a + (7 – 1)d] = 11[a + (11 – 1 )d] ⇒ 7(a + 6d) = 11(a + 10d) ⇒ 7a + 42d = 11a + 110d ⇒ 4a = -68d ⇒ a = -17d ∴ a18 = a + (18 – 1)d = a + 17d = -17d + 17d = 0